ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

7. Lie Groups and Lie Algebras

1991 MS Classification: 22, 58, 11 Lizhen Ji, University of Michigan, Ann Arbor, USA The trace class conjecture in the Selberg trace formula An important tool in understanding automorphic representations is the Selberg trace formula, and an open problem in the theory of Selberg trace formula is the trace class conjecture which asserts that for a smooth test function of compact support, the restriction of the convolution operator to the discrete subspace is of the trace class. I will present a solution to this conjecture, which is also solved independently by W. Müller.